OpenStudy (babydoll.m):

ABCD is a rectangle. What is the value of x? rectangle with length 45 feet, width x feet, and diagonal 53 feet 28 feet 45 feet 53 feet 98 feet

OpenStudy (babydoll.m):

@BioHazard9064 plz help

OpenStudy (trojanpoem):

 \[(53)^{2} = 45^{2} + x^{2}\] \[x = \sqrt{ 53^2 - 45^2}\]

OpenStudy (trojanpoem):

x = 28 ft.

OpenStudy (babydoll.m):

thank u @trojanpoem would you mind helping with some more

OpenStudy (trojanpoem):

Nope, I wouldn't mind.

OpenStudy (babydoll.m):

The length of the hypotenuse of a right triangle is 24 inches. If the length of one leg is 8 inches, what is the approximate length of the other leg? 16.0 inches 18.4 inches 22.6 inches 25.3 inches

OpenStudy (babydoll.m):

@TrojanPoem

OpenStudy (trojanpoem):

 x = sqrt(24^2 - 8^2) =